OpenStudy (anonymous):

i will become a fan and i will give you a medal if you can help me with this question

OpenStudy (anonymous):

Felix exclaims that his quadratic with a discriminant of −1 has no real solutions. Felix then puts down his pencil and refuses to do any more work. Create an equation with a negative discriminant. Then explain to Felix, in calm and complete sentences, how to find the solutions, even though they are not real.

OpenStudy (blake57roger):

i dont know this awnser

OpenStudy (xmoses1):

Yeah i don't know this either

OpenStudy (anonymous):

@Preetha @whpalmer4 @Zedd @Dobby1 @shamil98

OpenStudy (whpalmer4):

Do you know what the discriminant is?

OpenStudy (anonymous):

its a formula right

OpenStudy (whpalmer4):

If you have a quadratic: \[ax^2+bx+c =0\]The discriminant is \(b^2-4ac\), which you may recognize as the thing under the radical sign in the formula for the solutions of a quadratic: \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

If we have a "perfect square", such as \((x+2)^2 = x^2+4x+4\), then the discriminant is equal to 0, and our solution is \[x = -\frac{b}{2a}\]You might even recognize that as the formula for the x-coordinate of the vertex of a parabola.

OpenStudy (whpalmer4):

If we don't have a perfect square, and the discriminant is positive, then our solutions look like\[x = \frac{-b\pm d}{2a}\]where \(d\) is a positive number.

OpenStudy (anonymous):

nope not really

OpenStudy (whpalmer4):

If we graph the equation, those values of \(x\) will be the points at which the parabola crosses the x-axis, also known as the zeros or the roots or the solutions.

OpenStudy (whpalmer4):

And the final case is where the discriminant is negative. What happens when you take the square root of a negative number?

OpenStudy (anonymous):

ok well the discriminate number is negative 1 and it dont have a perfect square

OpenStudy (anonymous):

so like what do we do

OpenStudy (whpalmer4):

I asked you a question: what happens when you take the square root of a negative number?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

how on looking through my notes now

OpenStudy (anonymous):

yea i cant seem to find it

OpenStudy (whpalmer4):

Well, let's do some thinking. What is the square root of 4?

OpenStudy (anonymous):

2

OpenStudy (whpalmer4):

Okay, and we know that is true because 2*2 = 4. What is the square root of -4?

OpenStudy (anonymous):

-2

OpenStudy (whpalmer4):

I thought you might say that. What does\[-2*-2 =\]

OpenStudy (anonymous):

oh it would be a postive 4

OpenStudy (whpalmer4):

Right! What kind of a number can we multiply by itself and get a negative number? A positive number * a positive number gives us a positive number. A negative number * a negative number also gives us a positive number.

OpenStudy (anonymous):

a negative and a positive

OpenStudy (whpalmer4):

Ah, but then it won't be multiplying by itself, will it?

OpenStudy (anonymous):

oh yea...i can think of anything

OpenStudy (anonymous):

*cant

OpenStudy (whpalmer4):

It's like we need to have some sort of number we can multiply by itself and get -1. Then we could factor the square root of -4 like this: \[\sqrt{-4} = \sqrt{-1*4} = \sqrt{-1}*\sqrt{4} = 2\sqrt{-1}\]We could give this special number a name, \(i\), and use it like a variable when we do algebra, but always remembering that \[i=\sqrt{-1}\]\[i^2=-1\]

OpenStudy (whpalmer4):

\[\sqrt{-4} = \sqrt{4}*\sqrt{-1} = 2i\] Going the other way: \[2i*2i = 4i^2\]but by definition, \(i^2=-1\) so that becomes \[2i*2i = 4i^2 = 4*-1 = -4\]And that's exactly what we want from the square root of -4, right?

OpenStudy (anonymous):

right

OpenStudy (whpalmer4):

Going back to our problem, if we have a negative discriminant \((b^2-4ac < 0)\), that means we'll have a square root of a negative number in our solutions to the quadratic equation. A square root of a negative number is an imaginary number (which is why we chose \(i\)). An imaginary number combined with a real number is a complex number, which can be written in the form \(a+bi\) where \(a,b\) are real numbers and \(i = \sqrt{-1}\)

OpenStudy (whpalmer4):

Shall we do an example of solving a quadratic that has imaginary or complex solutions?

OpenStudy (anonymous):

yes please

OpenStudy (whpalmer4):

Let's pick one where \[a=1,\,b=2,\,c=3\] \[y=1x^2+2x+3\]\[y=x^2+2x+3\] What will the value of the discriminant be?

OpenStudy (anonymous):

now wouldnt we use the quadratic formula

OpenStudy (whpalmer4):

please, when I ask you a question, do me the courtesy of answering it...

OpenStudy (anonymous):

\[\sqrt[-2]{-8} \over 4\]

OpenStudy (whpalmer4):

No, I'm not sure what that is supposed to be, but it isn't the discriminant :-) If your quadratic is in the form \[ax^2+bx+c=0\]the discriminant is \[\Delta =b^2-4ac\](that's a Greek letter called delta, and it is often used as a symbol for the discriminant) What is the value of the discriminant for \[x^2+2x+3=0\]\[\Delta=\]?

OpenStudy (anonymous):

it would be -8

OpenStudy (whpalmer4):

very good!

OpenStudy (anonymous):

yayyyy

OpenStudy (whpalmer4):

So let's proceed on to find the solutions. Can you fill in the blanks in the quadratic formula? \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a} = \frac{-b\pm\sqrt{\Delta}}{2a}\]

OpenStudy (whpalmer4):

actually, before you do that, let's just focus on the \(\sqrt{\Delta}\) part. You found that \(\Delta=-8\), so we need to find the value of \[\sqrt{-8}\]Can you simplify that at all?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

and i had already put in the quadratic equation

OpenStudy (whpalmer4):

Can you simplify \[\sqrt{8}\]?

OpenStudy (anonymous):

4

OpenStudy (whpalmer4):

Does 4*4 = 8?

OpenStudy (anonymous):

1*8=8

OpenStudy (whpalmer4):

What's your point?

OpenStudy (whpalmer4):

Are those the only numbers which you can multiply to get 8?

OpenStudy (anonymous):

2*4

OpenStudy (whpalmer4):

Right. Can you take the square root of either of those numbers?

OpenStudy (whpalmer4):

\[\sqrt{8} = \sqrt{2*4} = \sqrt{2}*\sqrt{4} = \]

OpenStudy (anonymous):

4 which is 2

OpenStudy (whpalmer4):

So that simplifies to...

OpenStudy (anonymous):

2

OpenStudy (whpalmer4):

No, \[\sqrt{8} \ne 2\]\[\sqrt{8} = \sqrt{2}*\sqrt{4} = \]

OpenStudy (whpalmer4):

\[\sqrt{8}=\sqrt{2}*\sqrt{4}=\sqrt{2}*2=2\sqrt{2}\]

OpenStudy (whpalmer4):

But we have \[\sqrt{-8}\]How do we form the square root of a negative number?

OpenStudy (whpalmer4):

Remember that example I did, where I showed that \[\sqrt{-4} = 2i\]? What does that suggest to you?

OpenStudy (anonymous):

oh yea i had forgot about that

OpenStudy (whpalmer4):

okay, well, reread that section of the book we're writing here and tell me what the answer is...

OpenStudy (anonymous):

i dont have a book

OpenStudy (whpalmer4):

"book we're writing here" — scroll back and read the example I mentioned...

OpenStudy (anonymous):

\[\sqrt{-8}=4i\]

OpenStudy (whpalmer4):

:-( If that's true, then \[4i*4i=-8\]\[4i*4i=4*4*i*i = 16i^2 = 16(-1) = -16\]

OpenStudy (anonymous):

ughhhhhh

OpenStudy (whpalmer4):

Look, \[\sqrt{-a} = \sqrt{-1*a} = \sqrt{-1}*\sqrt{a} = i\sqrt{a}\]So if \[\sqrt{8} = 2\sqrt{2}, \sqrt{-8} =\]

OpenStudy (anonymous):

-2 sqrt 2

OpenStudy (whpalmer4):

I think you should go to this web site: https://www.khanacademy.org/math/arithmetic/exponents-radicals watch the videos in the section labeled "The Square Root"

OpenStudy (whpalmer4):

No, that's not correct. \[-2\sqrt{2}*-2\sqrt{2} = -2*-2*\sqrt{2}*\sqrt{2} = 4*2 = 8\]

OpenStudy (whpalmer4):

Again, if the square root of a positive number is \(a\), then the square root of the negative number will be \(ai\). square root of 8 is \(2\sqrt{2}\) square root of -8 is \(2\sqrt{2}i \text{ or } 2i\sqrt{2}\)